6 The school administration would ke to know who is taking the city bus to school A survey of 200 students showed that the buses held 98 freshman, 65 sophomores, 16 juniors, and 21 seniors. Use this information to make the ollowing a. A vertical bar graph b. A Pareto chant A circle graph Senior
Answer #1

(a)

Vertical bar graph based on the data for all categories of students is shown below.

Bar Chart of Students travelling taking city bus Freshman Sophomores Juniors Seniors

(b) Pareto Chart -

The frequency, cumulative frequency, percentage and cumulative percentage are given in tabular format as,

Frequency Cumulative Frequency Percentage Cumulative Percentage
Freshman 98 98 98/200 = 49% 49%
Sophomores 65 98+65 = 163 65/200 = 32.5% 49+32.5 = 81.5%
Junior 21 163 + 21 = 184 21/200 = 10.5% 81.5+10.5 = 92%
Senior 16 184 + 16 = 200 16/200 = 8% 92 + 8 = 100%

The pareto chart, based on cumulative frequency and cumulative percentage is shown below.

Pareto Chart of x 200 100% 80% 150 60% 100 40% 50 20% 0% び)

c) Circle graph

The frequency and percentage are given in tabular format as,

Frequency Percentage
Freshman 98 98/200 = 49%
Sophomores 65 65/200 = 32%
Junior 21 21/200 = 10%
Senior 16 16/200 = 8%
Total 200 100%

Circle graph based on the precentages is,

Pie Chart of Students travelling taking city bus Freshman 49% Seniors 10% Sophomores 32% Juniors 8%